For the Years Ended December 31, (in millions) Operating activities Net carnings Noncash items 2017 2016 2015 $10,823 7,073 S 5,868 Depreciation and amortization Deferred income taxes Share-based compensation Other, net 2,245 2,055 ,693 (965) 597 217 (73) 406 (82) 235) 81 485 Net change in other operating items, net of effects from acquisitions and changes in AARP balances: Accounts receivable Other assets Medical costs payablo Accounts payable and other liabilities Unearned revenues ,062) 357(591) (630) 1,601) (1,430) 1,284 ,849 2,585 930 1,494 ,280 157 (202) 237 13,596 9,795 9,740 Cash flows from operating activities Investing activities Purchases of investments Sales of investments Maturities of investments Cash paid for acquisitions, net of cash assumed Purchases of property, equipment and capitalized software Other, net Cash flows used for investing activities (14,388) (17,547) (9,939) 4,623 7,339 6,054 5,646 4,2813,354 (2,131) (1,760) (16,164) (2,023) (1,705) ,556) 37 (144) (8,599) 9,355) (18,395) (126)

Explain and describe what the statement is showing. For example, what do the values mean? Is it good? Is it bad?
Answer #1

The statement is showing the Cash Flow of the operating and investing activities. The operating activities shows a positive cash flow to organization of $9740m, $9795m and $13596m in the year 2015, 2016 and 2017 respectively. The increase in the cash flow from year 2015 to year 2017 shows a positive growth.

Whereas, the investing activities is showing negative cash flow through the years, which means that net cash has been applied or invested in assets. The cash flow outside has been reducing from $18395m to $9355m and to $8599 in the year 2015, 2016 and 2017 respectively. It is showing that organization has been reducing its investment and getting in positive cash flow positive.

In totality, the net cash flow has been getting better as :

Year 2015 = 9740m – 18395m = -8655m

Year 2016 = 9795m – 9355m = 440m

Year 2017 = 13596m – 8599m = 4997m

The cash flow has improved over the years. It shows that in the initial years of business, the organization has to face the negative cash flow, which get positive in future as the investment generates results. The table shows a very positive and good cash flow.